# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 2.3 Finite Automaton States
Thus the state set
$S(M,T)=$ $\{\varphi_{ij}(t_{i}),\varphi_{I}(t_{j}),\varphi_{O}(t_{k}),\varphi_{S}(t_{j}),\varphi_{A}(t_{j}),$ (2.8)
$i=1,\ldots,n,j=1,\ldots,n-1,k=2,\ldots,n\}$
describes $M$’s inputs, outputs, and state-transitions over $T$. Here $Cn(X)$ denotes the logical closure of the set of well-formed formulas $X$. This completes the proof.
It follows that every finite automaton can be formalized by the state set given in Theorem 2, which moreover exhibits atuomatic transtion behavior. Hence we introduce :
Definition 4 (Finite Automaton State): If a state set $S(X,T)$ expresses a finite automaton, then $S(X,T)$ is called a finite automaton state.
2.4 Definition of Information
As shown inXu2024 , the necessary and sufficient conditions for satisfying Postulates 1-4 are the existence of the following non-empty sets: an ontological set $o$, a set of times for the occurrence of the ontology $T_{h}$, a set of ontological states $S_{o}(o,T_{h})$, a set of objective carriers $c$, a set of times for the reflection of the carrier $T_{m}$, and a set of carrier reflection states $S_{c}(c,T_{m})$, such that $S_{o}(o,T_{h})$ enables a mapping to $S_{c}(c,T_{m})$. Denote this mapping as $I$, or more simply, $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, which is called the sextuple model of information.
This leads to the general definition of information:
Definition 5 (Mathematical Definition of Information): Let $\mathbb{O}$ denote the set of objective entities, $\mathbb{S}$ the set of subjective entities, and $\mathbb{T}$ the set of times, with elements of $\mathbb{O}$, $\mathbb{S}$, and $\mathbb{T}$ partitioned as needed depending on the domain. The ontology $o\in\mathbb{O}\cup\mathbb{S}$, time of occurrence $T_{h}\subseteq\mathbb{T}$, ontological state set $S_{o}(o,T_{h})$ satisfying Postulate 5, carrier $c\subseteq\mathbb{O}$ , reflection time $T_{m}\subseteq\mathbb{T}$, and carrier reflection set $S_{c}(c,T_{m})$ satisfying Postulate 5 are all non-empty sets. Then, information $I$ is the enabling mapping from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$.The set of all information $I$ is called the information space, denoted by $\mathfrak{I}$, and is one of the three major constituents of the objective world.
It is important to note that both $S_{o}(o,T_{h})$ and $S_{c}(c,T_{m})$ are state representations that satisfy Postulate 5. According to Theorem 1, each can be interpreted as a set of well-formed formulas within some formal system. However, the corresponding formal systems for the ontology and the carrier need not be identical. This distinction reflects one of the essential characteristics of information: its ability to establish connections between disparate entities in the world.
Definition 5 does not require that the ontological states of information be mapped one-to-one onto the carrier states. Of course, in practical scenarios, one-to-one mappings often do exist, which can simplify the analysis considerably.
Definition 6 (Recoverable Information): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ still denote a bijection from $S_{o}(o,T_{h})$ to $S_{c}(c,T_{m})$, i.e., $I$ is invertible. Thus, there exists an inverse mapping $I^{-1}$ such that $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$. In this case, we say that information $I$ is recoverable, and $S_{o}(o,T_{h})$ is the recoverable state set of information $I$.
In particular, owing to the importance of information in information systems, we propose:
Definition 7 (Information in an Information System): Let $I=(o,T_{h},S_{o},c,T_{m},S_{c})$ be a recoverable information. If $S_{o}(o,T_{h})$ is a finite automaton state, then $c$ is called an information system, and $I$ is called information in the information system $c$.
For recoverable information, we also use $I^{-1}(S_{c}(c,T_{m}))\underline{\triangleleft}S_{o}(o,T_{h})$ or $I^{-1}:S_{c}(c,T_{m})\leftleftarrows S_{o}(o,T_{h})$ to maintain consistency in the direction of enabling. Since, as recoverable information, the carrier losslessly carries the complete state of the ontology, this naturally represents an ideal scenario in information studies. If $I$ is not a bijection, we can always partition elements of $S_{o}(o,T_{h})$ according to the same image equivalence classes, and then establish an enabling mapping from these equivalence classes to $S_{c}(c,T_{m})$ by means of equivalence class reduction, thereby ensuring recoverability. Specifically:
Proposition 1 (Recoverable Reduction of Information): For $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $R$ be an equivalence relation on $S_{o}(o,T_{h})$ defined by $xRy\leftrightarrow I(x)=I(y)$. Then, there exists a unique enabling mapping $\hat{I}:S_{o}(o,T_{h})/R\rightrightarrows S_{c}(c,T_{m})$, and $\hat{I}$ is a bijection, i.e., $\hat{I}$ is recoverable. In this case, $\hat{I}$ is called the recoverable reduction of information $I$.
When $I$ is already recoverable, $S_{o}(o,T_{h})/R=S_{o}(o,T_{h})$, and $\hat{I}=I$. Recoverable information and the recoverable reduction of information can provide significant assistance in simplifying problems.
Reference Shannon1948 is generated as a foundational work in information theory. The paper proposed that $H=-\sum p_{i}\log p_{i}$ plays a central role as the measure of information in information theory." The discussion in this paper includes both the source and channel of information. The source is the origin of the information, and it generates events $e_{1},e_{2},\ldots,e_{n}$ with probabilities $p_{1},p_{2},\ldots,p_{n}$, respectively. The channel $I_{c}$ is the medium through which information is transmitted. After an event $e_{i}$ occurs at time $T_{h}$ in the source $I_{s}$, the channel $I_{c}$ receives and transmits the corresponding code $c_{i}$ in the time interval $(T_{h},T_{h}+\tau)$(where $1\leq i\leq n,\tau>0$).We define the ternary predicate $IsEvent^{3}(I_{s},T_{h},e_{i})$ to represent that event $e_{i}$ happened in source $I_{s}$ at time $T_{h}$, and the ternary predicate $IsCode^{3}(I_{c},T_{m},c_{i})$ to represent that the channel $I_{c}$ received and transmitted code $c_{i}$ in time interval $T_{m}$. Then,
$S_{I_{s}}(I_{s},T_{h})=IsEvent^{3}(I_{s},T_{h},e_{i})$ (2.9)
and
$S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))=IsCode^{3}(I_{c},(T_{h},T_{h}+\tau),c_{i})$ (2.10)
are the states of the source $I_{s}$ at time $T_{b}$ and the state of the channel $I_{c}$ in the time interval $(T_{h},T_{h}+\tau)$, respectively (where $1\leq i\leq n)$. Furthermore,
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.11)
represents the functional enabling mapping from the source state $S_{I_{s}}(I_{s},T_{h})$ to the channel state $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$.From equation (2.10), we can recover the event $e_{i}$ in equation (2.9) using the code $c_{i}$, thus $I$ is still recoverable information. Therefore, although Shannon’s information theory investigates the various probability distributions of $S_{I_{s}}(I_{s},T_{h})$ and the complex implementation mechanisms of $S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$, it forms a rich and extensive theoretical system. The OIT definition of information can encompass Shannon’s most original model of information, indicating that it has a broader and more universal application range.
2.5 Transmissibility of Information
Xu2014 ; xu2015 discuss transmissibility using composite mappings. Building on Definitions 1 and 5, we can characterize the fundamentals of information transmission more rigorously:
Theorem 3 (Transmissibility of Information) If $I_{1}=\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\rangle$ is information from an ontology $o$ to a carrier $c_{1}$, and $I_{2}=\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is information from carrier $c_{1}$ to carrier $c_{2}$, then the composition $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is also information, which we call transmission of information from $I_{1}$ via carrier $c_{1}$ to carrier $c_{2}$.
Proof: According to Definition 1, Definition 5, and the given conditions, for any $S_{o}(x,t)\in S_{o}(o,T_{h})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Furthermore, there exists $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$. We define $I_{3}(S_{o}(x,t))=S_{c_{2}}(x_{2},t_{2})$. Since the physical realization of $S_{c_{2}}(x_{2},t_{2})$ depends on the existence of $S_{c_{1}}(x_{1},t_{1})$, and the existence of $S_{c_{1}}(x_{1},t_{1})$ in turn depends on $S_{o}(x,t)$, it follows by transitivity of causality that the realization of $S_{c_{2}}(x_{2},t_{2})$ ultimately depends on $S_{o}(x,t)$.
Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{1}(S_{o}(x,t))\underline{\triangleright}S_{c_{1}}(x_{1},t_{1})$. Thus, $I_{3}(S_{o}(x,t))=I_{2}(I_{1}(S_{o}(x,t)))\underline{\triangleright}I_{2}(S_{c_{1}}(x_{1},t_{1}))\underline{\triangleright}S_{c_{2}}(x_{2},t_{2})$.
Therefore, $I_{3}=\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 5. In essence, this is the information $I_{1}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$. The theorem is thus proved.
Transmission is an important way of information flow. Communication, processing, input, and output are all concrete manifestations of information transmission. In many cases, information undergoes more complex flows through chain-like transmission mechanisms. From Theorem 3, we can readily derive the following:
Corollary 1 (Information Chain) Let $\{I_{i}=\langle o,T_{h},S_{o},c_{i},T_{m_{i}},S_{c_{i}}\rangle|i=1,\ldots,n\}$ be a sequence of information mappings, where for any $i>1$, it holds that $c_{i-1}=o$, $T_{h}=T_{m_{i-1}}$, and $S_{o}(o,T_{h})=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any $1\leq i\leq j\leq n$, $I_{ij}=\langle o,T_{h},S_{o},c_{j},T_{m_{j}},S_{c_{j}}\rangle$ also qualifies as information according to Definition 5.
2.6 Noisy Information
Many researchers question the significance of recoverable information, as in the real world, information is often not perfectly recoverable. This is because the reflection of information is often influenced by noise or other external factors. Therefore, the following concept is introduced:
Definition 8 (Noisy Information): For information $I=(o,T_{h},S_{o},c,T_{m},S_{c})$, let $\widetilde{S_{o}}(o,T_{h})$ be another state of the same carrier $c$ at time $T_{m}$ and $\widetilde{S_{c}}(c,T_{m})$ be another state of the same carrier $c$ at time $T_{m}$,
[…]
where the state $S^{-}_{o}(o,T_{h})$ is the loss state of $S_{o}(o,T_{h})$, satisfying $S^{-}_{o}(o,T_{h})\subseteq$ $S_{o}(o,T_{h})$ and not being enabled to map to $\widetilde{S_{c}}(c,T_{m})$, and the state $S^{+}_{o}(o,T_{h})$ is the superposed state of $S_{o}(o,T_{h})$, satisfying $S_{o}(o,T_{h})\cap S_{o}^{+}(o,T_{h})=\emptyset$ and being enabled to map to $\widetilde{S_{c}}(c,T_{m})$. The enabling mapping is then defined as:
$\tilde{I}:\widetilde{S_{o}}(o,T_{h})\rightrightarrows\widetilde{S_{c}}(c,T_{m})$ (2.13)
which is the noisy information of $I$, while $I$ is its own noiseless information. Thus, the noisy information $\tilde{I}$ is also information. It has the same ontology, time of occurrence, carrier, and reflection time as $I$, but the states of the entity and carrier may overlap but are not completely the same. Further analysis of equation (2.12) shows that information $I$ and $\tilde{I}$ are completely asymmetric, and hence:
Proposition 2 (Symmetry of Noisy Information): If $\tilde{I}$ is the noisy information of $I$, then $I$ is also the noisy information of $\tilde{I}$.
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information IIcannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenarioShannon1948 . In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
$E=f(S,N)$ (2.14)
where $f$ is the distortion function of the input for the transmitted signal. According to OIT’s view, the signal $S$ here is the ontological state of information $I$, $S_{I_{s}}(I_{s},T_{h})$. The distortion function $f(S,N)$ for signal $S$ and noise $N$ is equivalent to removing a state $S_{I_{s}}^{-}(I_{s},T_{h})$ from the source state $S_{I_{s}}(I_{s},T_{h})$ and adding another state $S_{I_{s}}^{+}(I_{s},T_{h})$. The output $E$ is then the channel’s output state $\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$, where $(T_{h},T_{h}+\tau)$ is the time interval of the channel output. Thus,
$\tilde{I}:\widetilde{S_{I_{s}}}(I_{s},T_{h})\rightrightarrows\widetilde{S_{I_{c}}}(I_{c},(T_{h},T_{h}+\tau))$ (2.15)
is the noisy information of
$I:S_{I_{s}}(I_{s},T_{h})\rightrightarrows S_{I_{c}}(I_{c},(T_{h},T_{h}+\tau))$ (2.16)
[…]
@@ 2.6 Noisy Information
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels inShannon1948 , demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
@@ 3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of finite automaton state and information models established in Section 2, the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
@@ 3.1 Construction of Initial Information
Suppose the model under study is denoted as M, and let $T=\{t_{i}|i=1,\ldots,n\}$ be the set of time points at which M, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of M is essentially the realization of information
$I_{o}=\langle M_{D},T_{o_{h}},S_{M_{D}},M,t_{1},S_{M}\rangle$ (3.1)
@@
where the ontology $M_{D}$ of $I_{o}$ is the designer’s conceptualization of M; $T_{o_{h}}$ is the time at which the designer conceives M, and $S_{M_{D}}(M_{D},T_{o_{h}})$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier M of $I_{o}$ is the model itself; the reflection time $t_{1}$ is when M is constructed according to $S_{M_{D}}(M_{D},T_{o_{h}})$; and $S_{M}(M,t_{1})$ is the formalization of the actual physical state of M upon its completion. Thus, $\sup(T_{o_{h}})<\sup(t_{1})$.
@@
According to Theorem 2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
$\varphi_{o_{1}}=State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ (3.2)
@@
denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},T_{o_{h}})$ as envisioned by the designer are assigned to $M_{D}$.
$\varphi_{o_{2}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land State^{3}(M_{D},t_{2},\Phi^{2}(M_{D},t_{2}))$ (3.3)
@@
expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},T_{o_{h}})$, and the input carrier state $S_{M_{I}}(M_{I},t_{1})$ is provided and satisfies the learnability condition determined by the binary predicate $Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, $M_{D}$ can optimize its internal structures, parameters, and algorithmsto $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, assigning this to $M_{D}$. This demonstrates the model’s capacity for learning and optimization, where $M_{I}$ denotes the input device of M.
$\varphi_{o_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{2}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land Output^{3}(M,t_{2},S_{M_{O}}(M_{O},t_{2})))$ (3.4)
@@
indicates that, at time $t_{1}$, if $S_{M_{I}}(M_{I},t_{1})$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate $Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))$, then $M_{D}$ can process $S_{M_{I}}(M_{I},t_{1})$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$ to yield the output carrier state $S_{M_{O}}(M_{O},t_{2})$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of M.
@@
$S_{M_{D}}(M_{D},T_{o_{h}})=\{\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}\}$ (3.5)
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},T_{o_{h}})$, and $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ all represent sets, $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
$S_{M_{D}}(M_{D},T_{o_{h}})$ may represent the designer’s idea, a design draft on paper, or an electronic document. Through the information $I_{O}$, these are enabled and mapped into the physically operational state of the model M:
$S_{M}(M,t_{1})=\{\psi_{o_{1}},\psi_{o_{2}},\psi_{o_{3}}\}$ (3.6)
where $\psi_{o_{1}}\underline{\triangleleft}I_{o}(\varphi_{o_{1}})$, $\psi_{o_{2}}\underline{\triangleleft}I_{o}(\varphi_{o_{2}})$, and $\psi_{o_{3}}\underline{\triangleleft}I_{o}(\varphi_{o_{3}})$ are the enabling mappings of $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$, respectively. This process establishes M as a typical finite automaton and an operational information system.

### v10

@@ 3.2 Noisy Information
In this way, we use the definition of noisy information to explain Shannon's model of noisy channels in (Shannon, 1948), demonstrating that OIT is applicable to the complex scenarios of information research under noisy conditions.
3.3 Information Chain
Reference (Xu et al., 2014) discusses the transmissibility of information through composite mappings. Based on Definitions 1 and 2, the fundamental laws governing information transmission can be characterized more rigorously as follows:
Lemma 1 (Transmissibility of Information): Let $I_{0}=\left\langle o,T_{h},S_{o},c_{1},T_{m_{1}},S_{c_{1}}\right\rangle$ and $I_{1}=\left\langle c_{1},T_{m_{1}},S_{c_{1}},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ be two pieces of information. Then, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is also information. We refer to this as the information $I_{0}$ being transmitted from carrier $c_{1}$ to carrier $c_{2}$ via $c_{1}$.
Proof: According to Definitions 1 and 2, as well as the given conditions, for any $S_{o}(x,t)\in S_{o}\left(o,T_{h}\right)$, there exists $S_{c_{1}}\left(x_{1},t_{1}\right)\in S_{c_{1}}\left(c_{1},T_{m_{1}}\right)$such that $I_{0}\left(S_{o}(x,t)\right)\trianglerighteq S_{c_{1}}\left(x_{1},t_{1}\right)$. Furthermore, there exists $S_{c_{2}}\left(x_{2},t_{2}\right)\in S_{c_{2}}\left(c_{2},T_{m_{2}}\right)$ such that $I_{1}\left(S_{c_{1}}\left(x_{1},t_{1}\right)\right)\trianglerighteq S_{c_{2}}\left(x_{2},t_{2}\right)$. Define $I_{2}\left(S_{o}(x,t)\right)=S_{c_{2}}\left(x_{2},t_{2}\right)$. Since the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ is contingent upon the existence of $S_{c_{1}}\left(x_{1},t_{1}\right)$, and the physical realization of $S_{c_{1}}\left(x_{1},t_{1}\right)$ in turn depends on $S_{o}(x,t)$, by causal induction, the physical realization of $S_{c_{2}}\left(x_{2},t_{2}\right)$ ultimately depends on $S_{o}(x,t)$.
Conversely, for any $S_{c_{2}}(x_{2},t_{2})\in S_{c_{2}}(c_{2},T_{m_{2}})$, there exists $S_{c_{1}}(x_{1},t_{1})\in S_{c_{1}}(c_{1},T_{m_{1}})$ such that $I_{1}(S_{c_{1}}(x_{1},t_{1}))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$; and there also exists $S_{o}(x,t)\in S_{o}(o,T_{h})$ such that $I_{0}(S_{o}(x,t))\trianglerighteq S_{c_{1}}(x_{1},t_{1})$. Thus, $I_{2}(S_{o}(x,t))={I_{1}(I}_{0}(S_{o}(x,t)))\trianglerighteq I_{1}(S_{c_{1}}\left(x_{1},t_{1}\right))\trianglerighteq S_{c_{2}}(x_{2},t_{2})$.
Therefore, $I_{2}=\left\langle o,T_{h},S_{o},c_{2},T_{m_{2}},S_{c_{2}}\right\rangle$ is an enabling mapping from $S_{o}(o,T_{h})$ to $S_{c_{2}}(c_{2},T_{m_{2}})$, which satisfies the definition of information as given in Definition 2. In essence, this is the information $I_{0}$ transmitted from carrier $c_{1}$ to carrier $c_{2}$. The lemma is thus proved.
Corollary 2 (Information Chain) Let $\left\{I_{i}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{i},T_{m_{i}},S_{c_{i}}\right\rangle\middle|i=1,\cdots,n\right\}$ be a sequence of information mappings, where for any $i>1$, it holds that $o_{i}=c_{i-1}$, $T_{h_{i}}=T_{m_{i-1}}$, and $S_{o_{i}}\left(o_{i},T_{h_{i}}\right)=S_{c_{i-1}}(c_{i-1},T_{m_{i-1}})$. In this case, we say that these mappings constitute an information chain. Moreover, for any 1 $\leq i\leq j\leq n$, $I_{ij}=\left\langle o_{i},T_{h_{i}},S_{o_{i}},c_{j},T_{m_{j}},S_{c_{j}}\right\rangle$ also qualifies as information according to Definition 2.
Transmission constitutes a fundamental mode of information flow and application. Communication, processing, input, and output are all concrete realizations of information transmission. In many cases, information undergoes more complex flows through chain-like transmission processes. Fig.1(c) intuitively illustrates such a process. For any pair of states within this process, the latter can be regarded as the result of the enabling mapping from the former, and this enabling mapping can be used to reasonably explain the causation of the latter state. This provides a theoretical foundation for the interpretability of the model.
@@ 4 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of information model and finite automaton state (Appendix B), the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
@@ 4.1 Construction of Initial Information
Suppose the model under study is denoted as $M$, and let $T=\left\{t_{i}\middle|i=1,\ldots,n\right\}$ be the set of time points at which $M$, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of $M$ is essentially the realization of information
$Io=\left\langle M_{D},{To}_{h},S_{M_{D}},M,t_{1},S_{M}\right\rangle$ (4.1)
@@
where the ontology $M_{D}$ of $Io$ is the designer’s conceptualization of $M$, ${To}_{h}$ is the time at which the designer conceives $M$, and $S_{M_{D}}\left(M_{D},{To}_{h}\right)$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier $M$ of $Io$ is the model itself; the reflection time $t_{1}$ is when $M$ is constructed according to $S_{M_{D}}\left(M_{D},{To}_{h}\right)$; and $S_{M}\left(M,t_{1}\right)$ is the formalization of the actual physical state of $M$ upon its completion. Thus, $\sup\left({To}_{h}\right)<sup(t_{1})$.
@@
According to Lemma B.2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
${\varphi o}_{1}={State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ (4.2)
@@
denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},{To}_{h})$ as envisioned by the designer are assigned to $M_{D}$.
$\begin{split}{\varphi o}_{2}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({Eq}^{2}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land{State}^{3}(M_{D},t_{2},\Phi^{2}\left(M_{D},t_{2}\right))\end{split}$ (4.3)
@@
expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},{To}_{h})$, and the input carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is provided and satisfies the learnability condition determined by the binary predicate ${Learnable}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, $M_{D}$ can optimize its internal structures, parameters, and algorithms to $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}\left(M_{I},t_{1}\right))$, assigning this to $M_{D}$. This demonstrates the model's capacity for learning and optimization, where $M_{I}$ denotes the input device of $M$.
$\begin{split}{\varphi o}_{3}=&{State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)\land{Input}^{3}(M_{D},t_{1},S_{M_{I}}\left(M_{I},t_{1}\right))\\ &\land{Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\\ &({{Eq}^{2}\left(S_{M_{O}}\left(M_{O},t_{2}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))\right)\land Output}^{3}(M,t_{2},S_{M_{O}}\left(M_{O},t_{2}\right)))\end{split}$ (4.4)
@@
indicates that, at time $t_{1}$, if $S_{M_{I}}\left(M_{I},t_{1}\right)$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate ${Processable}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$, then $M_{D}$ can process $S_{M_{I}}\left(M_{I},t_{1}\right)$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{1}\right))$ to yield the output carrier state $S_{M_{O}}\left(M_{O},t_{2}\right)$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of $M$.
@@
$S_{M_{D}}\left(M_{D},{To}_{h}\right)=\left\{{\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}\right\}$ (4.5)
@@
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},{To}_{h})$, and ${State}^{3}\left(M_{D},t_{1},\Phi^{2}(M_{D},{To}_{h})\right)$ all represent sets, ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
$S_{M_{D}}\left(M_{D},{To}_{h}\right)$ may represent the designer’s idea, a design draft on paper, or an electronic document. Through the information $Io$, these are enabled and mapped into the physically operational state of the model $M$:
$S_{M}\left(M,t_{1}\right)=\left\{{\psi o}_{1},{\psi o}_{2},{\psi o}_{3}\right\}$ (4.6)
where ${\psi o}_{1}\trianglelefteq Io({\varphi o}_{1})$, ${\psi o}_{2}\trianglelefteq Io({\varphi o}_{2})$, and ${\psi o}_{3}\trianglelefteq Io({\varphi o}_{3})$ are the enabling mappings of ${\varphi o}_{1},{\varphi o}_{2},{\varphi o}_{3}$, respectively. This process establishes $M$ as a typical finite automaton and an operational information system.